A sloped ceiling has a run of 4 and a rise of 1. What are the measurements of all the angles?
Answer by vleith(2983)

one angle is 90.
the smallest angle will be acrtan(1/4) = 14.03
The last angle = 180-90-14.03 = 75.97
A
| 1
--------4----B find the inverse Tangent (tan^-1 or arctan) of the angle formed at vertex B.